# NuRF: Nudging the Particle Filter in Radiance Fields for Robot Visual Localization

### v1

@@ IV Methodology
IV-A Anchor Setting in Radiance Fields
@@ 8       $\xi\leftarrow[s[1],s[2],0,s[3],0,0]$;
9       $\boldsymbol{F}\leftarrow\boldsymbol{F}\cup$ Embedding($I$);
@@ IV-B Nudging by Visual Place Recognition
After generating the VPR anchors, we need to incorporate the reference information provided by the VPR results into the pixel-wised weighting particle filter to solve Problem 9. We adopt the definition of the nudging step provided by Akyildiz[18], illustrated in Definition IV.1. In this definition, $x_{\tau}\in\mathsf{X}$ represents the system state at time $\tau$, and $y_{\tau}$ denotes the observations at time $\tau$. The state after the nudging step is denoted as $x^{+}$.
@@ IV-C Adaptive Scheme in NuRF
Figure 5: Particles are nudged by retrieving and utilizing images corresponding to anchors.
We introduce an adaptive scheme within the particle filter framework to switch the global localization and pose tracking. Initially, in global localization, we lower the resolution of both the input and rendered images. This reduces localization accuracy slightly but expands the area where high similarity responses are detected, lessening sensitivity to small movements. As the process iterates and the particles begin to converge, we maintain the original input image quality and enhance the resolution of the rendered images to improve pose tracking accuracy. This adaptive strategy dynamically adjusts the resolution based on the discrepancy between nudged and original particles. This approach effectively addresses the issues of particle deficiency in global localization and aids in automatic recovery from kidnapping scenarios (Problem III.3).
@@
$w_{\tau,j}=\frac{\text{SSIM}(I_{\tau},L(\mathcal{G},K^{-},T_{\tau,j}))}{\sum_{i=1}^{N}\text{SSIM}(I_{\tau},L(\mathcal{G},K^{-},T_{\tau,i}))}$ (14)
@@ V Experiments and Evaluation
We validate the proposed NuRF in a motion capture room, which is a typical indoor environment with cluttered objects. The experiments are designed to assess the performance in terms of global localization and pose tracking, i.e., the convergence from scratch and local tracking accuracy. Additionally, within the NuRF framework, we conduct a thorough comparison of three widely used image similarity functions: SSIM, complex wavelet-SSIM (CW-SSIM)[23], and LPIPS. These functions calculate image similarity in different spaces and allow us to assess the impact of the nudging step on the approximation of the posterior distribution in the particle filter. All the experiments are conducted on a blimp robot [24].
@@ V-A1 Radiance Field in Indoor Environments
V-A2 Image Database for Retrieval
In Section IV, an additional assumption is made that the radiance field could provide valuable visual place retrieval information to guide the particle behavior. Once the 3D radiance field is constructed, we then generate images from any desired pose within the model. We establish anchor points, denoted as $a=[x,y,\omega]$, uniformly on a two-dimensional sub-manifold of SE(3). Images are rendered from these anchor points and encoded into vectors using the Vision Transformer technique. The resulting vector representations, derived from the anchor point images, are then stored as vectors for the next steps.
@@ V-A3 Implementation Details
The experiments are conducted with Pytorch-based implementation. The computing hardware comprises an Intel CPU i7-13700KF and a GeForce RTX4080 GPU. The input image size was set to $680\times 800$, with a render resolution of $64\times 80$ for global localization and a render resolution of $680\times 800$ for pose tracking. The thresholds for global localization and pose tracking switching were set to $\lambda^{+}=5$, while the threshold for kidnapping detection was set to $\lambda^{-}=15$.
@@ V-B Global Visual Localization
In Figure 7, the first column illustrates the utilization of 400 particles. The initial positions of these particles were obtained by uniformly perturbing the dimensions of the room, which has a height of $1.8m$, a length of $5m$, and a width of $4m$. For orientation, the yaw angles were uniformly initialized within the range of $[-180^{\circ},180^{\circ}]$, while the pitch and roll angles for all particles were set to $0^{\circ}$. Once the experiment begins, the motion model provides a 6-DoF (six degrees of freedom) motion matrix to update the pose of the particle swarm (depicted in the second to fourth columns of Figure 7). The resampling process also occurs in the 6-DoF space, meaning that the algorithm optimizes the particles in a complete 6-DoF state and performs only one update per image.
Furthermore, we conduct experiments involving three different similarity functions: SSIM, CW-SSIM, and LPIPS in Figure 8. These experiments demonstrate that the NuRF algorithm can enable the pixel-wise difference function to achieve a similar level of localization mean squared error (MSE) median as the feature-based difference functions. The results of these experiments are presented in Figure 9, which showcases 20 trials with the designed NuRF.
Figure 8: Global visual localization with different similarity evaluation functions in sub-manifold. First Row: From left to right are the observed images of the robot in frame 1, frame 25, and frame 50, respectively. Second to Fourth Row: The global localization performance of NuRF when using different similarity evaluation functions.
Figure 9(a) showcases the accuracy of global localization and the convergence efficiency achieved with different similarity functions. The results demonstrate the exceptional performance of the NuRF method in the global localization task, even using the pixel-wised similarity metric to calculate the importance factor of the particles. When considering both running time and localization accuracy, the combined comparison in Figure 9(b) demonstrates that the pixel-wise similarity function outperforms the feature-wise similarity function in terms of median running cost. This outcome highlights the capability of the NuRF framework to address the limitations associated with using the pixel-wise similarity function as the likelihood function, thereby achieving a more efficient and effective localization process.
(a) Translational Mean Square Error
(b) Convergence time $\times$ Translational Mean Square Error
Figure 9: (a) The performance during the global localization task is depicted through violin plots derived from 20 independent trials. (b) We also show the product of convergence time and mean square translation error across all configurations.
@@ V-C Pose Tracking in Radiance Fields
In this section, we evaluate the performance of the NuRF method in a pose tracking task using images captured from a blimp. We initiate the tracking process with the pose estimation obtained in Section V-B. We employ $200$ particles for the pose tracking task and initialize their positions by sampling from a normal distribution centered around the estimated position. The variances for the position dimensions are set to $[0.2,0.2,0.1]$ respectively. Similarly, we sample the yaw, pitch, and roll angles from normal distributions centered around the estimated rotation. The variances for the angles are set to $[5,1,1]$ respectively. Figure 10 illustrates the tracking process, where only one update is performed per image.
@@
To validate the effectiveness of the NuRF, we conduct experiments on 20 randomly selected trajectories consisting of 40 frames each. These experiments aimed to demonstrate that the performance of the NuRF algorithm is not dependent on a specific similarity evaluation function. In order to further assess its adaptability, we repeated the tests using different similarity functions.
The results of orientation estimation are showing in Figure 11, the NuRF showcases its proficiency in accurately tracking yaw angles with an error margin of $15^{\circ}$ or less, irrespective of the similarity evaluation function employed (see Figure 11(a)). This emphasizes its robustness and effectiveness in precisely estimating and tracking yaw angles, regardless of the specific similarity metric utilized.
(a) Yaw Error
(b) Pitch Error
(c) Roll Error
Figure 11: (a) The average rotation error in yaw of 20 trails at each frame. The shaded area shows one standard deviation above and below the mean error. (b) The average rotation error in pitch of 20 trails at each frame. (c) The average rotation error in roll of 20 trails at each frame.
In high-resolution images, the similarity function tends to better recognize floors and ceilings in indoor scenes. The experimental results presented in Figure 11(b) show that the NuRF algorithm achieves stable tracking of the pitch angle of the camera pose within an error margin of $10^{\circ}$, regardless of the specific similarity function utilized. However, the experimental results in Figure 11(c) indicate that the tracking performance of the roll angle is influenced by the choice of similarity function. In real-time scenarios, employing feature-wise similarity functions like CW-SSIM and LPIPS for heading angle estimation entails significant computational power consumption and introduces errors in the roll direction during the resampling process. However, the use of the pixel-wise similarity function, complemented by nudging information, enables more precise and stable tracking of the robot’s pose. This suggests that the NuRF algorithm effectively compensates for and mitigates these errors, resulting in reliable roll angle tracking.
@@ V-D Ablated Experiments
TABLE I: Ablated Experiments on Anchor Number for VPR
Similarity
function
Anchor
number
Average
translational
MSE($m$)
Average
convergent
steps
FPS($Hz$)
0 1.204 64.7 0.26
SSIM 504 0.951 60.1 0.23
2502 0.642 24.2 0.21
0 1.233 74.1 0.15
CW-SSIM 504 1.185 70.7 0.14
2502 0.741 35.3 0.12
0 0.984 50.3 0.07
LPIPS 504 0.865 50.9 0.07
2502 0.517 22.5 0.07
The first ablated experiment is designed to evaluate global localization capabilities under two specific conditions: without VPR nudging (where the anchor point count was set to zero) and with attenuated VPR nudging (where the number of nudging particles was halved and the anchor point count was reduced to 504).
The results show that if there are no anchored points, i.e., no nudging step is performed, there is a noticeable decline in the accuracy. Specifically, the translational mean square error of position estimation decreases from 0.642 $m$ to 1.204 $m$ when using SSIM for visual place recognition. In terms of efficiency, the experimental results demonstrate that the nudging step leads to a slight increase in the time required for a single particle update, approximately 1 second. However, it significantly reduces the number of iteration steps needed for the convergence.
Following the convergence of global localization, the second ablation experiment assesses the impact of image resolution on pose tracking accuracy. By adjusting the resolution of rendered images, we explore its effect on the precision of the pose tracking algorithm. Figures 12, 13, and 14 present a comparison of estimation errors across different Euler angles using various rendering resolutions and perception functions. The results clearly show that using high-resolution images typically results in smaller estimation errors in two-thirds of the cases. Furthermore, a comparison with Figures 11 reveals that low-resolution images tend to perform better in the relatively simpler task of estimating the pitch angle, but they are 83% more likely to yield poorer results in the more challenging yaw and roll angle estimations. Overall, these findings strongly indicate that high-resolution images generally enhance pose tracking performance across the system.
(a) SSIM
(b) CW-SSIM
(c) LPIPS
Figure 12: (a) The comparison of average rotation error in yaw of 20 trails at each frame when using SSIM in different resolution. (b) The comparison of average rotation error in yaw of 20 trails at each frame when using CW-SSIM in different resolutions. (c) The comparison of average rotation error in yaw of 20 trails at each frame when using LPIPS in different resolutions.
(a) SSIM
(b) CW-SSIM
(c) LPIPS
Figure 13: (a) The comparison of average rotation error in the pitch of 20 trails at each frame when using SSIM in different resolutions. (b) The comparison of average rotation error in the pitch of 20 trails at each frame when using CW-SSIM in different resolutions. (c) The comparison of average rotation error in the pitch of 20 trails at each frame when using LPIPS in different resolutions.
(a) SSIM
(b) CW-SSIM
(c) LPIPS
Figure 14: (a) The comparison of average rotation error in the roll of 20 trails at each frame when using SSIM in different resolutions. (b) The comparison of average rotation error in the roll of 20 trails at each frame when using CW-SSIM in different resolutions. (c) The comparison of average rotation error in the roll of 20 trails at each frame when using LPIPS in different resolutions.
@@ VI Conclusions and Future Study
We consider several promising directions that can be explored to advance NuRF. One potential direction is to investigate methods for determining the minimum number of anchors required to achieve satisfactory localization performance. Another promising direction is the integration of planning capabilities within the radiance fields, thus achieving full navigation with such implicit representations.

### v2

@@ IV Methodology
To address the outlined problems, we introduce the NuRF framework as depicted in Figure 2. This framework integrates anchor points in close proximity to the target image $I_{\tau}$ within the feature space during the resampling phase. By implementing this strategy, both Problem 6 and Problem III.1 are effectively addressed for real-world deployment. Additionally, the utilization of nudging particles facilitates the monitoring of particle dispersion, i.e., the variance of particles, which is used to switch the global localization and poser tracking in the radiance fields.
IV-A Radiance Fields-enhanced VPR Anchors
Figure 3: We show the original image and the rendered image in pixel space and feature space correspondingly. In the comparison column, the difference between the original image and the rendered image is highlighted in blue. We assess the similarity between a query image and images rendered in the same orientation but shifted along the X-axis using pixel-wise similarity and feature-wise similarity.
@@ 8       $\xi\leftarrow[s[1],s[2],0,s[3],0,0]$;
9       $\boldsymbol{F}\leftarrow\boldsymbol{F}\cup$ Encode($I$);
@@ IV-B Nudging by Visual Place Recognition
After generating the VPR anchors, we need to incorporate the reference information provided by the VPR results into the pixel-wised weighting particle filter to solve Problem 6. We adopt the definition of the nudging step provided by Akyildiz[23], illustrated in Definition IV.1. In this definition, $x_{\tau}\in\mathsf{X}$ represents the system state at time $\tau$, and $y_{\tau}$ denotes the observations at time $\tau$. The state after the nudging step is denoted as $x^{+}$.
@@ IV-C Adaptive Scheme in NuRF
Figure 5: Pipeline of nudging particles using the feature information from VPR. Upon the observation of a new image, the ViT encoder transforms it into a feature map. This map is then vectorized and subjected to cosine distance calculations with the vectors stored in the VPR database. The camera pose corresponding to the $k$ vectors with the shortest cosine distance are retrieved and add into the particle set.
We introduce an adaptive scheme within the particle filter framework to switch the global localization and pose tracking. Initially, in global localization, we lower the resolution of both the input and rendered images. This reduces localization accuracy slightly but expands the area where high similarity responses are detected, lessening sensitivity to small movements. As the process iterates and the particles begin to converge, we maintain the original input image quality and enhance the resolution of the rendered images to improve pose tracking accuracy. This adaptive strategy dynamically adjusts the resolution based on the discrepancy between nudged and original particles. This approach effectively addresses the issues of particle deficiency in global localization and aids in automatic recovery from kidnapping scenarios.
@@
$w_{\tau,j}=\frac{\text{LOSS}(I_{\tau},L(T_{\tau,j}))}{\sum_{i=1}^{N}\text{LOSS}(I_{\tau},L(T_{\tau,i}))}$ (11)
@@ V Experiments and Evaluation
We validate the proposed NuRF in a motion capture room, which is a typical indoor environment with cluttered objects. The experiments are designed to assess the performance in terms of global localization and pose tracking, i.e., the convergence from scratch and local tracking accuracy. All the experiments are conducted on a blimp robot [28].
@@ V-A1 Radiance Field in Indoor Environments
V-A2 Baseline
We compare NuRF with Loc-NeRF [10] in both global visual localization and pose tracking tasks to demonstrate its superior accuracy and speed. NuRF utilizes 3DGS as the map, which renders much faster than NeRF [8] in the original Loc-NeRF. Therefore, to ensure a fair comparison, we replace the NeRF with the 3DGS map in the Loc-NeRF system.
@@ V-A3 Implementation Details
The experiments are conducted with Pytorch-based implementation. The computing hardware comprises an Intel CPU i7-13700KF and a GeForce RTX4080 GPU. The input image size was set to $680\times 800$, with a render resolution of $64\times 80$ for global localization and a render resolution of $680\times 800$ for pose tracking. The thresholds for global localization and pose tracking switching were set to $\lambda^{+}=5$.
@@ V-B Global Visual Localization
A demonstration for global visual localization in radiance fields with NuRF is shown in Figure 7, the first column illustrates the utilization of 400 particles. The initial positions of these particles were obtained by uniformly perturbing the dimensions of the room, which has a height of $1.8m$, a length of $5m$, and a width of $4m$. For orientation, the yaw angles were uniformly initialized within the range of $[-180^{\circ},180^{\circ}]$, while the pitch and roll angles for all particles were set to $0^{\circ}$. Once the experiment begins, the motion model provides a 6-DoF (six degrees of freedom) motion matrix to update the pose of the particle swarm (depicted in the second to fourth columns of Figure 7). The resampling process also occurs in the 6-DoF space, meaning that the algorithm optimizes the particles in a complete 6-DoF state and performs only one update per image.
Figure 8: Average localization errors of NuRF and comparisons over 20 trials.
Figure 9: Average time costs of NuRF and comparisons over 20 trials.
We conducted a test to compare the global localization capabilities of the Loc-NeRF algorithm with ours using the same image sequence. Additionally, to validate the modules in NuRF, we performed ablation experiments by omitting the nudging step and utilizing only the bootstrap particle filter (BPF) with pixel-wise weights for global localization performance. The experimental results are displayed in Figures 8 and 9. In the box plot, the upper and lower edges indicate the maximum and minimum localization errors from the 20 experiments, while the box length represents the interquartile range. The median is denoted by a blue dashed line in the center of the box. Figure 8 shows that the median localization error for NuRF is approximately 0.5 m, which is lower than that of Loc-NeRF and the BPF when the VPR nudging step is removed. Additionally, NuRF exhibits the most stable localization error across the 20 experiments, with significantly lower variance compared to the other two algorithms. Figure 4 further highlights the operational efficiency of NuRF, illustrating that its convergence time is notably shorter than that of Loc-NeRF and the standard BPF without the nudging step.
@@ V-C Pose Tracking in Radiance Fields
In this section, we evaluate the performance of the NuRF method in a pose tracking task using images captured from a blimp. We initiate the tracking process with the pose estimation obtained in Section V-B. We employ $200$ particles for the pose tracking task and initialize their positions by sampling from a normal distribution centered around the estimated position. The variances for the position dimensions are set to $[0.2,0.2,0.1]$ respectively. Similarly, we sample the yaw, pitch, and roll angles from normal distributions centered around the estimated rotation. The variances for the angles are set to $[5,1,1]$ respectively. Figure 10 illustrates a case study of tracking process, where only one update is performed per image.
@@
To validate the effectiveness of the NuRF, we conduct experiments on 20 randomly selected trajectories consisting of 40 frames each and repeat the same experiment on Loc-NeRF and BPF with these data. The results of the experiment are presented in Table I, which also shows the mean attitude tracking error and position error for each algorithm after 40 time steps. It can be seen in Figure 11 that the algorithm NuRF has smaller tracking error than Loc-NeRF in pitch, yaw, and 3D spatial position.
(a) Translation
(b) Roll
(c) Pitch
(d) Yaw
Figure 11: Average translational and orientational error of the robot position over 40 frames of 20 trials.
TABLE I: Average tracking error of various algorithms for 20 trials.
Average orientation
tracking error (degree)
Average position
tracking error (m)
Roll Pitch Yaw
Loc-NeRF 3.51 12.93 55.91 1.72
NuRF (ours) 2.17 7.81 12.18 0.64
@@ V-D Ablated Experiments
The ablated experiment is designed to evaluate global localization capabilities under two specific conditions: without VPR nudging (which is degenerated to pixel-wised BPF) and with attenuated VPR nudging (where the number of nudging particles was halved and the VPR anchor point count was reduced to 504). The results show in Table II that if there are no anchored points, i.e., no nudging step is performed, there is a noticeable decline in the accuracy. Specifically, the translational mean square error of position estimation decreases from 0.642 $m$ to 1.204 $m$ when using NuRF for global localization. In terms of efficiency, the experimental results demonstrate that the nudging step leads to a slight increase in the time required for a single particle update, approximately 1 second. However, it significantly reduces the number of iteration steps needed for the convergence.
TABLE II: Result of ablated experiment for 20 trials.
Average position
tracking error (m)
Average run
frequncy (Hz)
Pixel-wised BPF 1.204 0.0909
504-NuRF 0.79 0.0769
2502-NuRF 0.64 0.0706
@@ VI Conclusions and Future Study
The proposed NuRF still has certain limitations for practical use. One key limitation is its deployment in texture-poor environments, where visual place recognition might fail due to the lack of sufficient visual features. Another limitation lies in the operational efficiency of NuRF. The main efficiency bottleneck is the rendering speed of images using NeRF. Currently, NuRF achieves a localization frequency of nearly 0.1 Hz. While this frequency is acceptable for low-speed ground mobile robots and indoor blimps, it poses challenges for integration into high-speed robotic platforms such as drones or self-driving cars.
We consider several promising directions to advance NuRF. One promising direction is the use of incremental 3D Gaussian Splatting [29], which could help adapt and refine the NuRF system for larger-scale environments. The use of Gaussian Splatting could also improve the operational efficiency compared to the original NeRF. Another direction involves integrating planning capabilities within radiance fields to enable full navigation. Additionally, future research will focus on engineering efforts to achieve real-time robot localization, which is crucial for deploying NuRF in high-speed robotic platforms and dynamic environments.